When talking a penalty kick, footballer applied a force of 30.0 N for a period of 0.05 s. If the mass of the ball is 0.075 kg, Calculate the speed with which the ball moves off.


1
Expert's answer
2022-01-17T09:59:05-0500

"F = ma"

"a = \frac{F}{m}= \frac{30}{0.075}=400\frac{m}{s^2}"

"v_0= 0"

"a = \frac{v_1-v_0}{t}"

"v_1 = at= 400*0.05=20\frac{m}{s}"

"\text{Answer: }20\frac{m}{s}"